OpenStudy (anonymous):

I WILL GIVE MEDAL 1)which gives the first two steps evaluate 3-2(7+1)²-1 a. Subtract 2 from 3,add 7 plus 1 b. add 7 plus 1, multiply 2 times 8 c. add 7 plus 1, square 8 d. Add 7 plus 1, subtract 1 form 8. 2-which of the following equalities are valid?Choose all that apply. a.(2+3)²-2=9 b.(2+3)² -2= 23 c.(6-2)²+6=8 d(6-2)² +6=22 3-Which statement illustrates the identity property of multiplication? a.9*1=9 b.9+1=10 c.9*0=0 d.9+0=9 4. Which of the following statements illustrates the identity property of addition? a.3X+1X= 4X b.3X+0=3X c.3X*1= 3X d.3X*0=0

OpenStudy (anonymous):

c. add 7 plus 1, square 8

OpenStudy (anonymous):

b.(2+3)² -2= 23

OpenStudy (anonymous):

a.9*1=9

OpenStudy (anonymous):

a.3X+1X= 4X